Plan
of the Day
- June 25-28, 2011
- June 24, 2011
- June 23, 2011
- June 22, 2011
- June 21, 2011
- June 20, 2011
- June 17-20, 2011
- June 16, 2011
- June 15, 2011
- June 14, 2011
- Weekend June 10-13, 2011
- June 8, 2011
- June 7, 2011
- Weekend June 4-6, 2011
- June 3, 2011
- June 2, 2011
- June 1, 2011
- Holiday Weekend May 27-31, 2011
- May 26, 2011
- May 25, 2011
- May 24, 2011
- Weekend
May 20-23, 2011
- May
19, 2011
- May
18, 2011
- May
17, 2011
- May
16, 2011
- Weekend
May 13-16, 2011
- May
12, 2011
- May
11, 2011
- May
10, 2011
- Weekend
May 6-9, 2011
- May
6, 2011
- May
5, 2011
- May
4, 2011
- May
3, 2011
- Weekend
Apr 29-May
2, 2011
- Apr
28, 2011
- Apr
27, 2011
- Apr
26, 2011
- Apr
25, 2011
- Weekend
Apr 22-24, 2011
- Apr
22, 2011
- Apr
21, 2011
- Apr
20, 2011
- Apr
19, 2011
June 25-28, 2011
Continue 1 hour stores; minimize turnaround time.
Injection intensity 1.7e9/bunch. Maximum Tandem/injector intensity for all fills.
STAR low intensity run, one 2 hour store.
Physics ends 2200 Tuesday 28 June.
APEX and Power Supply ramp tests follow.
Operations end at 0600 Wednesday 29 June.
Main ring hi-pot and cryo warm-up follows.June 24, 2011
Continue 1 hour stores; minimize turnaround time.
Injection intensity 1.7e9/bunch. Maximum Tandem/injector intensity for all fills.
If down time arises, access for:
- PHENIX A/C, in IR or berm
- PHENIX RPC, in tunnel
- Blue abort kicker module inspection
- HLRF
June 23, 2011
Continue 1 hour stores; minimize turnaround time.
Injection intensity 1.7e9/bunch. Maximum Tandem/injector intensity for all fills.June 22, 2011
Try for 27 GeV Physics. Start store lengths at 1-2 hours, at discretion of experiments at this point.
Take more drift correction data between stores, adjust functions if necessary.
Continue tuning up ATR efficiency.June 21, 2011
Continue 27 GeV setup.
- Setup ramp collimation
- Revisit store collimation if necessary
- Measure drift at injection between ramps
- Put one bunch in both rings immediately after reaching injection
- turn BPM aquisition on at 4sec event, if not on already.
- turn on BBQ and check lock.
- add some more bunches to blue and yellow if needed for better bbq signal
- take 15-20 min data
Return to optimal Tandem setup at first available opportunity after NSRL completes their day.
Start Physics stores.June 20, 2011
APEX 0800-1600
27 GeV setup follows. Possible experimenter access prior to beam development.
- switch to ramp Au11v6
- 3+ hysteresis ramps.
- Re-establish injection
- ramp with feedback
- store setup
- measure and correct for drift at injection
- measure and correct for drift at store if need be
June 17-20, 2011
Continue Physics. Maximum bunch intensity.June 16, 2011
Dump store at 0900. You may want to arrange prior stores accordingly
Machine Development starts at 0900:
- LLRF firmware upgrades, up to 4 hours.
- Last test ramp for LLRF is first ramp for AnDY crossing angle/collisions/ZDC calibration work:
- remove IP2 crossing angle
- remove separation bumps, find collisions
- repeat on normal ramp, 2 4-hour stores
- Make ramp chromaticity measurement at some point if time allows.
Return to Physics after 2 stores with AnDY in collision.June 15, 2011
Maintenance 0800-1600
Recover PhysicsJune 14, 2011
Continue 4 hour Physics stores.
Intensity goal 1.6e9/bunch. - Keep pushing Tandem and injector intensities.
- Keep tuning x line quads to recover Blue injection efficiency. Document your work.
- Have best beam ready for fill. See RHIC cycle
checklist.
One store with a few low intensity bunch collisions:
- Find a Tandem pulse width setting that gives you 25% nominal intensity in RHIC.
- load emitTest fill pattern in RhicInjection
- To fill:
- Take one AGS cycle, nominal intensity
- change chopper
- Take one AGS cycle, low intensity
- restore chopper
- Fill ring.
- Switch rings and repeat 1-5 above
- ramp
Possible MD: remove crossing angle at IP2.Weekend June 10-13, 2011
Continue 4 hour Physics stores.
Intensity goal 1.6e9/bunch. - Keep pushing Tandem and injector intensities.
- Keep tuning x line quads to recover Blue injection efficiency. Document your work.
- Have best beam ready for fill. See RHIC cycle
checklist.
June 8, 2011
Test store: no cooling for comparison.
At
present beam intensities, we would like to see the difference cooling
makes. Ideally, we should make 2 stores as identical as possible,
the only difference being the lack of stochastic cooling.
- Follow a good store with cooling in all planes
- Turn off cooling in ssop
- Ramp with identical intensity to previous ramp
- Keep for at least 3 hours, then dump by experiment consensus.
Continue 4 hour Physics stores.
Intensity goal 1.6e9/bunch. - Keep pushing Tandem and injector intensities.
- Tune x line quads to recover Blue injection efficiency
- Have best beam ready for fill. See RHIC cycle
checklist.
June 7, 2011
Continue 4 hour Physics stores.
Intensity goal 1.6e9/bunch. - Keep pushing Tandem and injector intensities.
- Tune x line quads to recover Blue injection efficiency
- Have best beam ready for fill. See RHIC cycle
checklist.
Weekend June 4-6, 2011
Continue 4 hour Physics stores.
Intensity goal 1.6e9/bunch. - Keep pushing Tandem and injector intensities.
- Tune x line quads to recover Blue injection efficiency
- Have best beam ready for fill. See RHIC cycle
checklist.
June 3, 2011
Continue 4 hour Physics stores.
Vernier scans, early in ramp. Thursday or Friday stores.
Additional access to repair BL cooling drive.
Intensity goal 1.6e9/bunch. - Keep pushing Tandem and injector intensities.
- Tune x line quads to recover Blue injection efficiency
- Have best beam ready for fill. See RHIC cycle
checklist.
June 2, 2011
Continue 4 hour Physics stores.
Intensity goal 1.6e9/bunch. - Keep pushing Tandem and injector intensities.
- Tune x line quads to recover Blue injection efficiency
- Have best beam ready for fill. See RHIC cycle
checklist.
Machine Development during day, end of physics store: beta* squeeze 2 IRs all planes.
Machine Development during day, 4 hours starting between stores: working point scan.
Vernier scans, early in ramp. Thursday or Friday stores.
Post-maintenance, or Thursday MD test ramp activity: no rebucketing, calibrate storage cavity
voltage YS4, YS5, BS4, BS5. Brennan/Blaskiwiecz/HLRF group.
June 1, 2011
Continue 4 hour Physics stores.
Intensity goal 1.6e9/bunch. - Keep pushing Tandem and injector intensities.
- Tune x line quads to recover Blue injection efficiency
- Have best beam ready for fill. See RHIC cycle
checklist.
Maintenance Day 0600-1600.
- Includes STAR, maybe PHENIX polarity reversals.
Other upcoming items:
- Vernier scans, early in ramp. Thursday or Friday stores.
- Post-maintenance, or Thursday MD test ramp activity: no rebucketing, calibrate storage cavity
voltage YS4, YS5, BS4, BS5. Brennan/Blaskiwiecz/HLRF group.
- Thursday MD activity: working point scan.
Holiday Weekend May 27-31, 2011
Continue 4 hour Physics stores.
Intensity goal 1.6e9/bunch. - Keep pushing Tandem and injector intensities.
- Tune x line quads to recover Blue injection efficiency
- Have best beam ready for fill. See RHIC cycle
checklist.
Not yet scheduled:
- Test ramp activity: no rebucketing, calibrate storage cavity
voltage YS4, YS5, BS4, BS5. Brennan/Blaskiwiecz/HLRF group.
- MD activity: working point scan.
May 26, 2011
Resume Physics stores, 4 hour length.
Intensity goal 1.5e9/bunch.
- Keep pushing Tandem and injector intensities.
- Tune x line quads to recover Blue injection efficiency
- Have best beam ready for fill. See RHIC cycle
checklist.
MD Thursday: working point scan with rebucketed beam
- starts between stores during day tomorrow.
Test ramp activity: no rebucketing, calibrate storage cavity
voltage YS4, YS5, BS4, BS5. Brennan/Blaskiwiecz/HLRF group.
Test ramp activity: replay ramp to gauge effecty of BBQ on ramp emittance growth.May 25, 2011
Continue Physics stores, 4 hour length -- to be revisited if cooling
gives significant lifetime improvements.
APEX 0800-2359
Intensity goal 1.5e9/bunch.
- Keep pushing Tandem and injector intensities.
- Tune x line quads to recover Blue injection efficiency
- Have best beam ready for fill. See RHIC cycle
checklist.
Test ramp activity: no rebucketing, calibrate storage cavity
voltage YS4, YS5, BS4, BS5. Brennan/Blaskiwiecz/HLRF group.
Test ramp activity: replay ramp to gauge effecty of BBQ on ramp emittance growth.
May 24, 2011
Continue Physics stores, 4 hour length -- to be revisited if cooling
gives significant lifetime improvements.
Intensity goal 1.5e9/bunch.
- Keep pushing Tandem and injector intensities.
- Have best beam ready for fill. See RHIC cycle
checklist.
End of store activity: beta squeeze tests.
Test ramp activity: no rebucketing, calibrate storage cavity
voltage YS4, YS5, BS4, BS5. Brennan/Blaskiwiecz/HLRF group.
Test ramp activity: replay ramp to gauge effecty of BBq on ramp emittance growth.
MD Tuesday afternoon: working point scan with rebucketed beamWeekend
May 20-23, 2011
Continue Physics stores, 4 hour length -- to be revisited if cooling
gives significant lifetime improvements.
Intensity goal 1.5e9/bunch.
- Keep pushing Tandem and injector intensities.
- Have best beam ready for fill. See RHIC cycle
checklist.
Stochastic cooling all planes on all stores if available.
End of store activity: check yellow collimation, y5 steering.
End of store activity: beta squeeze tests.
Test ramp activity: no rebucketing, calibrate storage cavity
voltage YS4, YS5, BS4, BS5. Brennan/Blaskiwiecz/HLRF group.
Test ramp activity: working point scan with rebucketed beam
May
19, 2011
Continue Physics stores, 4 hour length.
PHENIX 30 min access during the day for MuID shielding work.
- Use available down time for access if possible
- Schedule access between stores near end of day shift
otherwise.
Test ramp activity: calibrate storage cavity voltage YS4,
YS5, BS4, BS5. Brennan/Blaskiwiecz/HLRF group.
Continue ramp and store tuning:
- Stochastic cooling setup on all available stores.
- Continue increasing injected intensity, to 1.5/bunch
- Yellow lifetime tuning at store.
May
18, 2011
Continue Physics stores, 4 hour length.
Maintenance 0800-1600. Booster off at 0600.
Test ramp activity: calibrate storage cavity voltage YS4,
YS5, BS4, BS5. Brennan/Blaskiwiecz/HLRF group.
Continue ramp and store tuning:
- Stochastic cooling setup on all available stores.
- Continue increasing injected intensity, to 1.5/bunch
- Yellow lifetime tuning at store.
May
17, 2011
Continue Physics stores, 4 hour length.
Machine Development ~1200-1600:
- At store, measure chromaticity. Check working
point. Yellow in particular.
- Blue injection efficiency tuning.
- flatten closed orbit at kicker
- injection orbit 4-bump?
- lambertson drive?
Continue ramp and store tuning:
- Stochastic cooling setup on all available stores.
- Continue increasing injected intensity, to 1.5/bunch
- Yellow lifetime tuning at store.
- Provide more storage voltage; condition cavities when
downtime allows.
May
16, 2011
Continue Physics stores, 4 hour length.
Continue ramp and store tuning:
- Stochastic cooling setup on all available stores.
- Continue increasing injected intensity, to 1.5/bunch
- Continue Blue injection efficiency tuning.
- Yellow lifetime tuning at store.
- Provide more storage voltage; condition cavities when
downtime allows.
Offline investigations:
- Blue injection historical data.
- Longitudinal emittance growth on ramp. Find
intensity dependencies in bunch-by-bunch data.
- WCM jump ~2.5 sec after transition?
Weekend
May 13-16, 2011
Continue Physics stores. Store length 4 hours.
Continue ramp and store tuning:
- Stochastic cooling setup on all available stores.
- Continue increasing injected intensity, to 1.5/bunch
- Continue Blue injection efficiency tuning.
- Provide more storage voltage; condition cavities when
downtime allows.
May
12, 2011
Continue Physics stores.
Continue ramp and store tuning:
- Stochastic cooling setup on all available stores.
- Increase injected intensity to 1.4/bunch
- Continue yellow transition efficiency tuning
- Continue Blue injection efficiency tuning.
- Provide more storage voltage; condition cavities when
downtime allows.
May
11, 2011
APEX and Maintenance 0001-1600
Continue Physics stores.
Continue ramp and store tuning:
- Increase injected intensity to 1.4/bunch
- Continue yellow transition efficiency tuning
- Continue Blue injection efficiency tuning.
May
10, 2011
Troubleshoot YA2 cavity.
Continue Physics stores.
Continue ramp and store tuning:
- Increase injected intensity to 1.3/bunch
- Increase to 111 bunches
- Continue yellow transition efficiency tuning (lower chrom
in gammat4 stone in 0.5 steps)
Weekend
May 6-9, 2011
Physics stores, 6 hour length.
Continue ramp and store tuning:
- Stochastic cooling setup on all available ramps.
- Continue ramp transmission tuning -- chromaticity before
transition, orbit after transition, losses at 175sec, etc.
- Increase bunch number and intensity if conditions allow.
May
6, 2011
Overnight stores for experiments.
Continue fine tuning of ramp and store:
- Debunching at injection. Check cavity phasing?
- Continue ramp collimation setup.
- Transition losses -- chromaticity?
- Orbit after transition.
- Dispersive losses (mid-arc) late in ramp
- Check and adjust Yellow store chromaticity
- Check injection drift correction
May
5, 2011
Continue store setup, and fine tune ramp.
- Rebucketing
- Check store chromaticity
- Continue tuning for store lifetime.
- Stochastic cooling -- call experts for any store.
- Set up and run ramp collimation
- Store automation - cogging on ramp, rebucketing, lisa
optimization, periodic SGOF at store
- Debunching beam at injection. Check acceleration
cavity phases (HLRF)?
- Work on transition losses, orbit excursions after transition
Overnight stores for experiments. Owl shifts are on call,
provided there is a store before midnight.
May
4, 2011
Maintenance begins 0600.
Resume store setup ~1700
- Restore beam
- check cogging, abort gap alignment
- Check YA1 operation
- Start rebucket setup
- Stochastic cooling setup
- Landau cavity setup
- Work towards 111 bunches 1.3/bunch
- store tune scans
- measure chromaticity (Yellow)
- measure optics
May
3, 2011
Continue ramp development:
- Troubleshoot YA1 RF cavity
- Set up ramp collimation
- Test RhicInjection orbit feedback corrections
- Wiggles ramp with Booster off (around NSRL schedule)
- Ramp cogging & separation bump removal, with
feedback on
Then store setup:
- Increase bunch intensity/number up to 56
- Steer for collisions
- Gap cleaning
- Collimation
- Rebucketing?
- 10 Hz feedback
- Stochastic cooling setup
- Increase bunch intensity/number > 56
Weekend
Apr 29-May
2, 2011
- Continue 30 minute stores.
- Continue to increase injector intensities/efficiencies.
- Last set of vernier scans.
- Run ends 0800 Monday.
- APEX ends 1400 Monday.
- 100 GeV setup begins Monday evening.
Apr
28, 2011
- Continue 30 minute stores.
- Continue to increase injector intensities/efficiencies.
- PHENIX zero field run at 11:00?
- Vernier scans if desired.
Apr
27, 2011
- Continue 30 minute stores.
- Continue to increase injector intensities/efficiencies.
- Examine RHIC closed orbit, injection efficiency.
- Vernier scans if desired.
Apr
26, 2011
- Continue 30 minute stores.
- Continue to increase injector intensities/efficiencies.
Apr
25, 2011
- Continue 30 min stores.
- Prepare for possible 1-2 hours Machine Development:
- Fix yellow, blue closed orbit offset
- Fix corrector avg offset trim
- Adjust bend trim
- Engage feedback
- Re-optimize collisions/collimation
- Engage 10 Hz GOFB
- Vernier scans TBD.
Weekend
Apr 22-24, 2011
- Work on collision setup
- Improve bunched beam lifetime
- Optimize collision rates, backgrounds
- use 56-bunch store and uncog/recog if needed to discern
collisions vs. background
- Inject cogged, colliding, and collimated
- Get appropriate signals, target rates from experiments
- Work on store automation
- Gap cleaning
- leave collimators in after dump.
- Automatic periodic orbit correction
- Other outstanding items
- 10 Hz GOFB
- Push injector intensities
- Injection kicker timing
- Continue troubleshooting B3 RF as time allows
Apr
22, 2011
- Continue beam setup
- Establish sufficient lifetimes for multi-bunch injection
(goal ~15% beam decay or less)
- Remove separation bumps at IP6 and IP8, steer for
collisions
- Develop injection with colliding beams
Apr
21, 2011
- Repair Booster B3 RF cavity.
- PHENIX access around 14:40 for DOE tour; requires
Restricted Access and PHENIX magnet off before 13:50.
- Continue beam setup:
- Re-establish circulating and captured beam
- Check tunes are near design values .22, .23
- Investigate orbit issues -- average corrector strength,
highly loaded correctors, etc.
- Establish sufficient lifetimes for multi-bunch injection
(goal ~15% beam decay or less)
- Inject multiple bunches
- Cog beams
- Remove separation bumps at IP6 and IP8, steer for
collisions
- Tune up cogged beams
- Develop injection with colliding beams
Apr
20, 2011
Access to assess and repair RF mechanical tuner on 28MHz acceleration
cavities.
Reverse q7 polarities for IBS-suppression lattice, in preparation for
100 GeV ramps
Switch to Au11v3 ramp
Continue AGS setup:
- tune for intensity
- investigate flattop losses
- set up and test multibunch extraction
- check ATR magman settings for IBS-lattice match
Continue RHIC setup at nominal injection energy 9.8 GeV:
- circulating beam setup, with sufficient intensity and
lifetime for capture
- RF capture
- Instrumentation setup
- Intensity and lifetime tuning
- Fine tune injection kicker timing
- monitor possible yellow bunch-kicker timing drift
Apr
19, 2011
Current schedule estimates Heavy Ion beam operations starting on this
day.
- Switch AGS to Au immediately following end of pp Jet run.
- work on AGS synchro, AGS extraction, ATR transport during
maintenance period.
- Switch ramps, test, run hysteresis (at least 3)
- Injection/first turn setup
- Try first-turn SVD from commissioning version of
RhicOrbitDisplay
- injection kicker setup